The sum of the squares of two consecutive integers is 13. Find the integers

Answer by stanbon(75887)

The sum of the squares of two consecutive integers is 13. Find the integers
-------
1st: x
2nd: x+1
------
Equation:
x^2 + (x+1)^2 = 13
x^2 + x^2 + 2x+1 = 13
2x^2 + 2x -12 = 0
-----
x^2 + x - 6 = 0
------
Factor:
(x+3)(x-2) = 0
x = 2 or x = -3
---
If x=2
1st = x = 2
2nd = (x+1) = 3
----
If x = -3
1st + x = -3
2nd = (x+1) = -2
=====================
Cheers,
Stan H.